Agenda/Minutes on 2021/1/22
CRY+VIS Interface Meeting
Participants: Takahashi, Sato, Hirata, Ikeda, Miyo, Ushiba, Tomaru, Kimura, Aso
- Schedule for IX: Cabling for the payload is scheduled from 1/26. Connection is scheduled in the week of 2/8.
- Schedule for IY: Cooling of payload was started from 1/18.
- Schedule for EX: Connection of ETMX was scheduled in the week of 3/15.
- Corrugate tube fixing jig: Re-measurement at the beginning of March (3/1-3) in IYV is planed. Deformation of the bottom ring of top-chamber would be large.
- Commissioning in vacuum:
- Acceptance check: Cryo-payload in IX and Type-A tower in EY are on going.
- Jig for FC painting: Under manufacturing.
- PF actuator: by the end of January.
Participants: Takahashi, Sato, Hirata, Miyo, Aso, Iguchi, Ishizaki, Tanaka, Washimi, Terrence
- Installation in EYV.
- Going to the F3 stage.
- Blades for FR in the F3 was replaced. Checked the motion of FR.
- Measured the spectrum of BF LVDT.
- Mounted the F3 on the EQ stop. The BF was suspended by the F3.
- Prepared the cables between the BF and F3.
- The cabling work from BF to F3 was completed.
- Stepper motor driver in EYV
One channel of the stepper motor driver for GAS had failed. k-log
Week of the 18th of January
Hirata-san worked in the design of the thermistor holder, cabling, etc.
I disconnected SRM picomotor drivers to move them away from the assembly frame. I failed to report it in a klogin due time.
Miyo-kun, Hirata-san and I reviewed the IP transformation matrices from IP LVDTs to L, T and Y and from actuation in L, T and Y to real coil-magnet actuator. We found a few of inconsistencies in existing documents. See klog reports 15769 and 15776.
- We recovered the IP alignment:
See klog 15782.
- I updated the medm screen with these matrices. Diagonalization is pending.
- I moved the IP close to the setpoint using horizontal FRs.
- I turned the control system on by hand and works fine. (I tried using Guardian but it keeps using the old setpoint values despite I took a snapshot.)
- I moved BF keystone to around zero. (F1 keystone has been in midrange for a long time and we should not move it unless it's absolutely necessary.)
- We went to the tunnel and recovered the oplev readout. We had to move the suspension in yaw by a lot (mrad maybe) using the new F0-Y stepper motor.
- After recovery I measured TFs in the IP. They look fine, but in IP-L and IP-T the coherence is low from 150 mHz to 400 mHz. Terrence suggested to check coherence with ground motion and we cna also increase the amount of actuation.
- We finally got to add ballast mass on top of BF:
- The final result is that the keystone is now about 16 um above the goal. It requires -1,100 counts of coil-magnet actuation to reach it. After pump down we can use the FR to move it closer if we think it's suitable.
- The aligmment of the OSEMs didn't change much.
- IM TF measuements reveal that the system is healthy.
- Things left to do in BS suspension:
- In F0 adjust ballast mass, FR position and limit switch positions.
- Tune IP resonant frequency.
- Measure IP-LVDT sensitivity.
- Install oplev box (what is the status?)
- Install outer frame large beam? (what is the status?)
Plan for next week
- Alignment recovery after cabling (it seems they want to use it next week for injecting green light.)
- Preparation for ballas mass adjustment.
[Fabian] What's the plan for next week? It is something like this:
- Traverser cable change?
- Gluing wire in PR3?
Although I told you that I explain my document for Suspension Commissioning, I should write more information. here
- Compared the TFs of modified LVDT board with the original board. Reduction of 50dB in the TF was found.
- Actual effect should be measured using the BF damper in PR.
- Type-B remedy work is on going.
- Installation of Type-A is on going in EY.
- Procedure documents.
- Criteria in the acceptance test.
- BS control noise
The week from 1/25
- Takahashi: 1/25-29
- Sato: 1/25-29
- Hirata: 1/25-29
- Yano: 1/25-29
The week from 2/1
- Takahashi: 2/1-5
- Sato: 2/1-5
- Hirata: 2/1-5
- Yano: 2/1-5
- On 2020/1/29(Fri)